OpenStudy (math_bender_15):

Medals and fans rewarded Two mechanics worked on a car. The first mechanic worked for 15 hours, and the second mechanic worked for 5 hours. Together they charged a total of $1625 . What was the rate charged per hour by each mechanic if the sum of the two rates was $175 per hour?

OpenStudy (anonymous):

Is it this one?

OpenStudy (math_bender_15):

yes

OpenStudy (anonymous):

You can get someones attention by using the @ character. For example @math_bender_15

OpenStudy (math_bender_15):

Thanks I didn't know that.

OpenStudy (anonymous):

So you can look at this question as simultaneous equations. Let's call mechanic1 x and mechanic2 y.\[15x +5y = 1625\]\[x+y =175\]Does that make sense?

OpenStudy (math_bender_15):

A little bit, so would I add 15 and 5 together then subtract that from 1625?

OpenStudy (anonymous):

No. Treat x and y as separate variables. you could divide equation 1 by 5.\[15x + 5y = 1625\]divided by 5\[3x+y = 325\]

OpenStudy (anonymous):

Now we have \[3x+y = 325\] and \[x+y = 175\]If we subtract equation 2 from eqaution1 we get.\[2x = 150 \rightarrow x= 75\]

OpenStudy (anonymous):

Meaning \[y = 100\]

OpenStudy (math_bender_15):

so that means that the first mechanic will make $75 an hour for the 15 that he worked and the second mechanic will make $100 an hour for the 5 that he worked.

OpenStudy (anonymous):

yep.

OpenStudy (math_bender_15):

Thank you, it makes sense to me now.